OpenStudy (itz_sid):

Determine whether the series is absolutely convergent, conditionally convergent, or divergent.

OpenStudy (itz_sid):

\[\sum_{n=0}^{\infty}\frac{ (-1)^n }{ 9n+1 }\]

OpenStudy (solomonzelman):

Well, if we just had \(\color{black}{\displaystyle \frac{1}{9n+1} }\) (without the (-1)^n part) would it converge?

OpenStudy (itz_sid):

If you look at the series without putting it in absolute values, does the graph look like this?